# Exact Diagonalization, Matrix Product States and Conformal Perturbation Theory Study of a 3D Ising Fuzzy Sphere Model

### v1

@@ 5.1 Hamiltonian
In this section we study the same microscopic Hamiltonian as in the pioneering work of Ref. Zhu2023 . The Hamiltonian is specified explicitly in App. D. For the present discussion the precise microscopics are not important. In essence the Hamiltonian has two effects or ingredients: i) the charged electrons carrying a spin [^18] This spin should be thought of as pseudospin since it does not couple to the orbital magnetic field. one-half degree of freedom move on the surface of the sphere under the effect of a perpendicular orbital magnetic field. The filling of the spherical Landau level is one electron per angular momentum orbital and under the effect of the interactions the electrons form a Mott insulating state, which means that the charge gap is finite, while the charges form an integer quantum Hall state with liquid-like charge correlations, i.e. they do not form a crystal. ii) The purpose of the interactions involving the spin degree of freedom of the electrons is to implement the competition between Ising interactions among spatially nearby electrons wanting to align the spin degrees of freedom ferromagnetically along the spin ±$\pm z$ direction on the one hand, and the transverse field term which strives to polarize the spins in the transverse, $x$ direction. At a suitably chosen value of the ratio of the transverse field to the other interaction parameters, a quantum phase transition takes place between the two phases, believed to be described by a 3D Ising CFT, and the results of Ref. Zhu2023 illustrate this beautifully.
@@ 5.2 Spectrum at criticality: Observation of finite size effects
While Ref. Zhu2023 presented an impressive number of primaries and descendants of the 3D Ising CFT with rather high accuracy, it is curious that for example the reported scaling dimension of the lowest primary, the $\sigma$ field, while being relatively accurate for small electron numbers, is converging only slowly with increasing system size. This motivated us to perform a more detailed analysis of the finite size effects governing the spectrum at small electron numbers. Given that the numerical results are already quite close to the results from conformal bootstrap, it is plausible to assume that the remaining corrections can be described using the tools of conformal perturbation theory Lao:2023zis .
@@
In Fig. 8 we show the low-lying spectrum of the fuzzy sphere transverse field Ising model for parameters $V_{0}=4.75$ and $h=3.16$, the same values as used in most of Ref. Zhu2023 . This spectrum has been obtained using exact diagonalization on up to 18 electrons (c.f. App. E.1), while for selected low lying levels on up to 32 electrons we used an MPS implementation (c.f. App. E.2) to obtain the corresponding energies.
Ref. Zhu2023 used the known scaling dimension $\Delta_{T}=3$ of the conserved stress energy tensor $T^{\mu\nu}$ to remove the speed of light from the spectrum. We will show below that from a CPT point of view the lowest primary $\sigma$ is actually least affected by perturbations other than the $\varepsilon$ field, while the effect of the many other possible perturbing fields on the stress energy tensor energy level are basically not known. Therefore we use the CPT framework to determine the speed of light $v$ and $g_{\varepsilon}$ from the lowest two energy gaps from the vacuum to $\sigma$ and to $\partial\sigma$ based on the known conformal bootstrap values of $\Delta_{\sigma}$ and $f_{\sigma\sigma\varepsilon}$ and their extension to $\partial\sigma$. The details of this procedure will be given in Sec. 5.4 below. We then rescale the spectrum to set $v/R$ to one (see Eqs. (2),(12)) and plot the resulting energy gaps, see Fig. 8. Since we are working at fixed microscopic parameters $V_{0}$ and $h$ there is a residual value of $g_{\varepsilon}(N)$ left, which we did not yet compensate for in Fig. 8 (because the effect of $g_{\varepsilon}$ is not known precisely for all the levels shown).
@@
Looking closely at the $\mathbb{Z}_{2}=-1$, $\ell=4$ spectrum in Fig. 8, the state counting hints at another primary not much above $\sigma^{\mu\nu\rho\sigma}$. Available conformal bootstrap data contain no sign of such a primary. In this respect, it will be interesting to see the data coming from the extremal functional method from the recent conformal bootstrap study Chang:2024whx .
@@ 5.3 Conformal Pertubation Theory for the 3D Ising CFT on the Fuzzy Sphere
Figure 9: 3D Ising $\delta E$ pattern. The panels are organized according to families of fields in Ref. (Simmons-Duffin2017, , Tables 3,4,5,7). The $x$-axis labels the $\sigma$ and $\varepsilon$ fields and their lowest descendants up to level 3. The $y$-axis enumerates the perturbing fields in the corresponding families.
In order to devise a practical CPT scheme for the 3D Ising CFT on the fuzzy sphere it is useful to discuss the known operators which are allowed to contribute to the CPT. On the sphere, we can use as a perturbing operator any $\mathbb{Z}_{2}$-even primary $\mathcal{U}=\mathcal{U}^{(\ell)}$ with even spin $\ell$ and even spatial parity (in an appropriate rotationally invariant contraction if $\ell>0$, see (28)). Several families of such primaries have been reported in Ref. Simmons-Duffin2017 together with the OPE coefficients $f_{\sigma\sigma\mathcal{U}}$ and $f_{\varepsilon\varepsilon\mathcal{U}}$.
To first order, shifts of the $\sigma$ and $\varepsilon$ primary energy levels corrections are given by Eq. (30), and for their descendants by the same equation times the relative correction factors, which for the first few descendant levels were computed Lao:2023zis ; notebook (see App. B). The precise values of the couplings $g_{\cal V}$ appearing in this equation are unknown (they have to be determined from a fit), but their order of magnitude can be estimated by dimensional analysis as
@@
In this setup, we plot in Fig. 9 the expected shifts of the $\sigma$ and $\varepsilon$ primaries and the descendants energy levels from perturbations associated with operators from Simmons-Duffin2017 . The figure is organized into 4 blocks which are grouped as in Ref. Simmons-Duffin2017 . In each subplot the vertical axis labels perturbing fields in the family. On the horizontal axis we arrange the first few energy levels in the $\sigma$ and $\varepsilon$ family. The intensity of the colored boxes (red positive, blue negative) displays the size of the expected correction. The plot reveals quite some structure, implying that not all energy levels are affected alike across energy levels and perturbing fields.
For example, one observes that the $\sigma$ energy level is basically only affected by the presence of a $g_{\varepsilon}$ couplings, while all the other shown perturbations have an insignificant effect. (As discussed previously the coupling to the stress energy tensor $T$ can be absorbed in the effective speed of light, for all states.) The first descendant of $\sigma$ is only affected by $g_{\varepsilon}$ and by $g_{C}$, where $C$ is the lowest spin-4 $\mathbb{Z}_{2}$-even primary. The $C$ perturbation is analogous to the perturbation $T^{2}+\bar{T}^{2}$ discussed in Sec. 4 in the 2D case. Note that the leading $\mathbb{Z}_{2}$-even irrelevant scalar primary $\varepsilon^{\prime}$ has little effect $\sigma$ and $\partial\sigma$, however in contrast it more significantly affects (=$\partial^{2}_{(\ell=0)}\sigma$, (=$\partial^{3}_{(\ell=1)}\sigma$ and the $\varepsilon$ family. Another observation is that descendants at small/large spin are more strongly affected by perturbations with small/large spin respectively.
@@
This complex structure of expected shifts renders it delicate to justify a fit of a large collection of perturbed energy levels obtained in numerics based only on a small number of perturbing couplings. Since we are operating at a relatively small number of electrons $N$, i.e. at small $R$, we cannot simply assume that only the leading irrelevant couplings contribute, but would need to infer potentially many couplings from fits. Another difficulty is that most OPE coefficients beyond the ones already discussed are not available from the conformal bootstrap, so that we can’t predict the corrections on other energy levels beyond $\sigma$ and $\varepsilon$ levels and their lowest few descendants. Finally, it so happens that some states belonging to different conformal multiplets have numerically close energy levels, raising the question about the importance of off-diagonal matrix elements and the associated level-mixing effects, which will be discussed in detail in Sec. 7.
In the following we will proceed by analyzing the effect of changes of microscopic parameters $h$ and $V_{0}$ on the energy spectrum, and we will infer which CPT couplings vary with $h$ and $V_{0}$.
We will see that the value of $V_{0}=4.75$ Zhu2023 has the virtue of having only small contributions due to $g_{\varepsilon^{\prime}}$ and $g_{C}$. Changing $h$ on the other hand is directly linked to controlling $g_{\varepsilon}$, as expected. Moving $V_{0}$ away from the fine tuned value $V_{0}=4.75$ used in Ref. Zhu2023 , both $g_{\varepsilon^{\prime}}$ and $g_{C}$ start to grow with opposite signs. In general we are currently not able to obtain reliable fits for further CPT couplings, because of the above-mentioned difficulties.
@@ 5.4 Minimal CPT model: fitting $v/R$ and $g_{\varepsilon}$
[…]
@@
We can then easily solve these formulas for $v/R$ and $g_{\varepsilon}$ and therefore obtain estimates for $v/R$ and $g_{\varepsilon}$ as a function of the microscopic coupling $h$. Note that this procedure relies on the known values of $\Delta_{\sigma}$ and $f_{\sigma\sigma\varepsilon}$ for the 3D Ising CFT. The spectrum in Fig. 10 is then shown after multiplying the ED gaps with $R/v$ (here we use $R=\sqrt{N}$).
@@
Fig. 10 is organized for clarity of presentation into four panels, each of which contains a subset of the energy levels. The leftmost panel displays $\sigma$ and its first three descendant levels. The second panel does the same for the $\varepsilon$ field. In the third panel we show the stress energy tensor and its level-one descendants. Finally in the last panel we show a few more primaries, both scalar and spinning. The thick lines present the ED data after removal of the speed of light. Thin lines of the same color show the CPT predictions $\Delta_{\mathcal{O}}+g_{\varepsilon}f_{\mathcal{O}\mathcal{O}\varepsilon}$, based on the extracted $g_{\varepsilon}$ values and the OPE coefficient $f_{\mathcal{O}\mathcal{O}\varepsilon}$ (times a relative correction factor in case of a descendant) which is either known from the conformal bootstrap or determined later in Sec. 6. The shaded regions highlight the difference between the ED level and the CPT model prediction.
@@
We see that the simple CPT model based on a determination of $g_{\varepsilon}$ by tracking $\sigma$ and $\partial\sigma$ as a function of $h$ and the knowledge of the OPE coefficients yields a very good agreement with surprisingly many other ED levels not included in the fit. A few levels are however not well captured by this simple CPT model, for example $\partial^{2}\varepsilon(\ell=0)$ and also the stress energy tensor are somewhat shifted vertically. So while the $h$ dependence is well captured by a model involving $g_{\varepsilon}$ to first order, a few ED levels differ from this model, and more sophisticated CPT models need to developed. We see that the remaining discrepancies (the widths of the shaded regions) are essentially independent of $h$ on the scale of this plot, suggesting that they may be corrected by couplings of irrelevant perturbations, which are approximately constant in a neighborhood of the critical point. The natural next step would be to add a coupling to the leading irrelevant scalar primary $\varepsilon^{\prime}$ Lao:2023zis ; Voinea:2024ryq . However, after having tried this we observed that this does not seem to improve the fit at $V_{0}=4.75$ too much, because many ED levels in principle affected by a finite $g_{\varepsilon^{\prime}}$ are already well described without invoking $g_{\varepsilon^{\prime}}$. So this correction is by itself unable to remedy the spectrum discrepancies without introducing discrepancies for energy levels where the agreement is already very good.
@@
It’s interesting to compare the situation of the fitting difficulties that we have here with the original work Lao:2023zis which advocated using CPT for describing the ED spectra in the icosahedron model, and with the subsequent work Voinea:2024ryq which applied CPT in the context of a fuzzy sphere model, where no such difficulties were noted. The main difference is that the criterion for declaring success was lower in Lao:2023zis ; Voinea:2024ryq , which dealt with the models which were not finetuned, so that the initial deviations from CFT were quite significant, and the dramatic improvement of the agreement after using the CPT was easier to see. Here on the contrary we are dealing with the model which was already carefully tuned in Zhu2023 by choosing $V_{0}=4.75$, to minimize deviations from CFT. Our CPT analysis shows that to reduce deviation when it is significant is easier than to improve agreement further when it is already small. A likely explanation is that some other small effects are at play, beyond those which are captured by our analysis, like level mixings discussed in Sec. 7, or perhaps something else. Isolating these effects with certainty remains an interesting open problem.

### v2

@@ 5.1 Hamiltonian
In this section we study the same microscopic Hamiltonian as in the pioneering work of Ref. [1]. The Hamiltonian is specified explicitly in App. D. For the present discussion the precise microscopics are not important. In essence the Hamiltonian has two effects or ingredients: i) the charged electrons carrying a spin [^19] This spin should be thought of as pseudospin since it does not couple to the orbital magnetic field. one-half degree of freedom move on the surface of the sphere under the effect of a perpendicular orbital magnetic field. The filling of the spherical Landau level is one electron per angular momentum orbital and under the effect of the interactions the electrons form a Mott insulating state, which means that the charge gap is finite, while the charges form an integer quantum Hall state with liquid-like charge correlations, i.e. they do not form a crystal. ii) The purpose of the interactions involving the spin degree of freedom of the electrons is to implement the competition between Ising interactions among spatially nearby electrons wanting to align the spin degrees of freedom ferromagnetically along the spin $\pm z$ direction on the one hand, and the transverse field term which strives to polarize the spins in the transverse, $x$ direction. At a suitably chosen value of the ratio of the transverse field to the other interaction parameters, a quantum phase transition takes place between the two phases, believed to be described by a 3D Ising CFT, and the results of Ref. [1] illustrate this beautifully.
@@ 5.2 Spectrum at criticality: Observation of finite size effects
While Ref. [1] presented an impressive number of primaries and descendants of the 3D Ising CFT with rather high accuracy, it is curious that for example the reported scaling dimension of the lowest primary, the $\sigma$ field, while being relatively accurate for small electron numbers, is converging only slowly with increasing system size. This motivated us to perform a more detailed analysis of the finite size effects governing the spectrum at small electron numbers. Given that the numerical results are already quite close to the results from conformal bootstrap, it is plausible to assume that the remaining corrections can be described using the tools of conformal perturbation theory [29].
@@
In Fig. 8 we show the low-lying spectrum of the fuzzy sphere transverse field Ising model for parameters $V_{0}=4.75$ and $h=3.16$, the same values as used in most of Ref. [1]. This spectrum has been obtained using exact diagonalization on up to 18 electrons (c.f. App. E.1), while for selected low lying levels on up to 32 electrons we used an MPS implementation (c.f. App. E.2) to obtain the corresponding energies.
Ref. [1] used the known scaling dimension $\Delta_{T}=3$ of the conserved stress energy tensor $T^{\mu\nu}$ to remove the speed of light from the spectrum. We will show below that from a CPT point of view the lowest primary $\sigma$ is actually least affected by perturbations other than the $\varepsilon$ field, while the effect of the many other possible perturbing fields on the stress energy tensor energy level are basically not known. Therefore we use the CPT framework to determine the speed of light $v$ and $g_{\varepsilon}$ from the lowest two energy gaps from the vacuum to $\sigma$ and to $\partial\sigma$ based on the known conformal bootstrap values of $\Delta_{\sigma}$ and $f_{\sigma\sigma\varepsilon}$ and their extension to $\partial\sigma$. The details of this procedure will be given in Sec. 5.4 below. We then rescale the spectrum to set $v/R$ to one (see Eqs. (2),(12)) and plot the resulting energy gaps, see Fig. 8. Since we are working at fixed microscopic parameters $V_{0}$ and $h$ there is a residual value of $g_{\varepsilon}(N)$ left, which we did not yet compensate for in Fig. 8 (because the effect of $g_{\varepsilon}$ is not known precisely for all the levels shown).
@@
Looking closely at the $\mathbb{Z}_{2}=-1$, $\ell=4$ spectrum in Fig. 8, the state counting hints at another primary not much above $\sigma^{\mu\nu\rho\sigma}$. Available conformal bootstrap data contain no sign of such a primary. In this respect, it will be interesting to see the data coming from the extremal functional method from the recent conformal bootstrap study [45].
@@ 5.3 Conformal Pertubation Theory for the 3D Ising CFT on the Fuzzy Sphere
Figure 9: 3D Ising $\delta E$ pattern. The panels are organized according to families of fields in Ref. [13, Tables 3,4,5,7]. The $x$-axis labels the $\sigma$ and $\varepsilon$ fields and their lowest descendants up to level 3. The $y$-axis enumerates the perturbing fields in the corresponding families.
In order to devise a practical CPT scheme for the 3D Ising CFT on the fuzzy sphere it is useful to discuss the known operators which are allowed to contribute to the CPT. On the sphere, we can use as a perturbing operator any $\mathbb{Z}_{2}$-even primary $\mathcal{U}=\mathcal{U}^{(\ell)}$ with even spin $\ell$ and even spatial parity (in an appropriate rotationally invariant contraction if $\ell>0$, see (28)). Several families of such primaries have been reported in Ref. [13] together with the OPE coefficients $f_{\sigma\sigma\mathcal{U}}$ and $f_{\varepsilon\varepsilon\mathcal{U}}$.
To first order, shifts of the $\sigma$ and $\varepsilon$ primary energy levels corrections are given by Eq. (30), and for their descendants by the same equation times the relative correction factors, which for the first few descendant levels were computed [29, 34] (see App. B). The precise values of the couplings $g_{\cal V}$ appearing in this equation are unknown (they have to be determined from a fit), but their order of magnitude can be estimated by dimensional analysis as
@@
In this setup, we plot in Fig. 9 the expected shifts of the $\sigma$ and $\varepsilon$ primaries and the descendants energy levels from perturbations associated with operators from [13]. The figure is organized into 4 blocks which are grouped as in Ref. [13]. In each subplot the vertical axis labels perturbing fields in the family. On the horizontal axis we arrange the first few energy levels in the $\sigma$ and $\varepsilon$ family. The intensity of the colored boxes (red positive, blue negative) displays the size of the expected correction. The plot reveals quite some structure, implying that not all energy levels are affected alike across energy levels and perturbing fields.
For example, one observes that the $\sigma$ energy level is basically only affected by the presence of a $g_{\varepsilon}$ couplings, while all the other shown perturbations have an insignificant effect. (As discussed previously the coupling to the stress energy tensor $T$ can be absorbed in the effective speed of light, for all states.) The first descendant of $\sigma$ is only affected by $g_{\varepsilon}$ and by $g_{C}$, where $C$ is the lowest spin-4 $\mathbb{Z}_{2}$-even primary. The $C$ perturbation is analogous to the perturbation $T^{2}+\bar{T}^{2}$ discussed in Sec. 4 in the 2D case. Note that the leading $\mathbb{Z}_{2}$-even irrelevant scalar primary $\varepsilon^{\prime}$ has little effect $\sigma$ and $\partial\sigma$, however in contrast it more significantly affects $\partial^{2}_{(\ell=0)}\sigma$, $\partial^{3}_{(\ell=1)}\sigma$ and the $\varepsilon$ family. Another observation is that descendants at low spin tend to be mostly affected by low-spin perturbations, while high-spin descendants couple to a larger variety of fields.
@@
The complex structure of expected shifts in Fig. 9 shows that a fit of a large collection of perturbed energy levels using a small number of perturbing couplings may be delicate to justify. Since we will be operating at a relatively small number of electrons $N$, i.e. at small $R$, one could worry about the need to infer potentially many couplings from fits. Another difficulty is that most OPE coefficients beyond the ones already discussed are not available from the conformal bootstrap, so that we can’t predict the corrections on other energy levels beyond $\sigma$ and $\varepsilon$ levels and their lowest few descendants. Finally, it so happens that some states belonging to different conformal multiplets have numerically close energy levels, raising the question about the importance of off-diagonal matrix elements and the associated level-mixing effects.
Below we will be optimistic and will use first-order CPT with at most three couplings: the relevant and the two leading irrelevant perturbations $g_{\varepsilon}$, $g_{\varepsilon^{\prime}}$, $g_{C}$. [^20] In particular, we will omit the next irrelevant operator $T^{\prime}$, of spin 2 and of dimension $5.50915(44)$ [13], i.e. somewhat above $C$. In the future it would be interesting to include $T^{\prime}$ in the fit. We will see that this procedure works already rather well, although some small discrepancies do remain. Because of these discrepancies and due to the above-mentioned difficulties, we are currently not able to extend the procedure and obtain reliable fits for further CPT couplings. The level-mixing effects will be discussed in detail in Sec. 7.
Our focus will be to infer how CPT couplings $g_{\varepsilon}$, $g_{\varepsilon^{\prime}}$, $g_{C}$ vary with $h$ and $V_{0}$. We will see that changing $h$ is directly linked to controlling $g_{\varepsilon}$, as expected. On the other hand $V_{0}$ controls $g_{\varepsilon^{\prime}}$ and $g_{C}$. We will see that the fine tuned value of $V_{0}=4.75$ [1] corresponds to small values of $g_{\varepsilon^{\prime}}$ and $g_{C}$. Moving $V_{0}$ away from $V_{0}=4.75$, both $g_{\varepsilon^{\prime}}$ and $g_{C}$ start to grow with opposite signs.
@@ 5.4 Minimal CPT model: fitting $v/R$ and $g_{\varepsilon}$
As for the 2D Ising model studied before, it is natural to expect that the transverse field $h$ is driving the transition from the $\mathbb{Z}_{2}$ ordered phase at small $h$ to the paramagnetic, massive phase at large $h$. We therefore expect that $h$ controls the coupling $g_{\varepsilon}$ which affects the finite volume CFT spectrum away from the critical point. We think of the term $(h-h_{c})\int_{S^{2}}d\Omega\,\sigma^{x}(\Omega)$ in a microscopic Hamiltonian sense as having representation in the CFT sense in the immediate vicinity of the critical point as $g_{\mathds{1}}\mathds{1}+g_{\varepsilon}\varepsilon+g_{\mu\nu}T^{\mu\nu}+\ldots$. We expect that coefficients for the left out operators will be small. Of those retained only $g_{\varepsilon}$ matters, as the other two ($g_{\mathds{1}}$ and $g_{\mu\nu}$) can be absorbed into an overall energy shift and a correction to the speed of light.
@@
We can then easily solve these formulas for $v/R$ and $g_{\varepsilon}$ and therefore obtain estimates for $v/R$ and $g_{\varepsilon}$ as a function of the microscopic coupling hh. [^21] While we do not show $v$ and $g_{\varepsilon}$ as a function of $h$ for this fit, we note that $v$ varies vary little in the shown range of $h$, while $g_{\varepsilon}(h)$ is very close to a plot in the second row of Fig. 11 for a more sophisticated fit. In particular $g_{\varepsilon}$ crosses zero at $h\approx 3.14$. Note that this procedure relies on the known values of $\Delta_{\sigma}$ and $f_{\sigma\sigma\varepsilon}$ for the 3D Ising CFT. The spectrum in Fig. 10 is then shown after multiplying the ED gaps with $R/v$ (here we use $R=\sqrt{N}$).
@@
Fig. 10 is organized for clarity of presentation into four panels, each of which contains a subset of the energy levels. The leftmost panel displays $\sigma$ and its descendants up to level $4$. The second panel provides the same information for $\varepsilon$ and its descendants up to level $3$. In the third panel we show the stress energy tensor and its level-one descendants. Finally in the last panel we show a few more primaries, both scalar and spinning. The thick lines present the ED data after removal of the speed of light. Thin lines of the same color show the CPT predictions $\Delta_{\mathcal{O}}+g_{\varepsilon}f_{\mathcal{O}\mathcal{O}\varepsilon}$, based on the extracted $g_{\varepsilon}$ values and the OPE coefficient $f_{\mathcal{O}\mathcal{O}\varepsilon}$ (times a relative correction factor in case of a descendant) which is either known from the conformal bootstrap or determined later in Sec. 6.
@@
We see that the simple CPT model based on a determination of $g_{\varepsilon}$ by tracking $\sigma$ and $\partial\sigma$ as a function of $h$ and the knowledge of the OPE coefficients yields a very good agreement with surprisingly many other ED levels not included in the fit. A few levels are however not well captured by this simple CPT model. Such levels show in Fig. 10 as shaded bands, where the shading highlight the difference between the ED level and the CPT model prediction. So while the $h$ dependence is well captured by a model involving $g_{\varepsilon}$ to first order, a few ED levels differ from this model, and more sophisticated CPT models need to developed. We see that the remaining discrepancies (the widths of the shaded regions) are essentially independent of $h$ on the scale of this plot, suggesting that they may be corrected by couplings of irrelevant perturbations, which are approximately constant in a neighborhood of the critical point. The natural next step would be to add a coupling to the leading irrelevant scalar primary $\varepsilon^{\prime}$ [29, 31]. However, after having tried this we observed that this does not seem to improve the fit at $V_{0}=4.75$ too much, because many ED levels in principle affected by a finite $g_{\varepsilon^{\prime}}$ are already well described without invoking $g_{\varepsilon^{\prime}}$. So this correction is by itself unable to remedy the spectrum discrepancies without introducing discrepancies for energy levels where the agreement is already very good.
@@
It’s interesting to compare the situation of the fitting difficulties that we have here with the original work [29] which advocated using CPT for describing the ED spectra in the icosahedron model, and with the subsequent work [31] which applied CPT in the context of a fuzzy sphere model, where no such difficulties were noted. The main difference is that the criterion for declaring success was lower in [29, 31], which dealt with the models which were not finetuned, so that the initial deviations from CFT were quite significant, and the dramatic improvement of the agreement after using the CPT was easier to see. Here on the contrary we are dealing with the model which was already carefully tuned in [1] by choosing $V_{0}=4.75$, to minimize deviations from CFT. In the future, it would be extremely important to isolate which other effects could explain the remaining discrepancies at $V_{0}=4.75$ and further improve the fit. One possibility is level mixing, which will be discussed in Sec. 7.